Consider the following subscription behavior information from Genie.com, a web site that provides tools for constructing a family tree (ancestor search). Subscriptions cost $9.99 per month, but you are charged for the entire year at the time of purchase. There is a one-year minimum term when you sign up for the service. Once purchased, subscriptions are set to renew automatically unless the subscriber cancels them. When a membership renews, it renews for a one-year term and again you are charged for the entire year. There are no variable costs associated with providing this service to an individual customer, but Genie does engage in customer relationship activities that they believe will increase customer retention. These customer relationship activities cost Genie about $10 per year per customer. Based on a sample of 1000 customers that joined Genie.com five years ago, near the time when the company was founded, they were able to determine how many of those customers remained subscribers in the second year, third year etc. Based on this information, Genie calculated the average annual retention rate to be 20%. Genie uses an annual discount rate of 8%.

a. Last year, Genie spent $10,000 placing advertisements on Google. Genie management believes that these advertisements were responsible for about 300 new subscribers. Would you recommend to Genie management that they purchase more Google ads?

b. Suppose a newly-introduced loyalty program increases the number of customers that remained to 30%. Does this new data change your answer to 9.a?

c. Do you have any hesitations or concerns about making recommendations to management based on your above estimate of customer lifetime value?
Answer #1
  1. Last year, Genie spent $10,000 placing advertisements on Google. Genie management believes that these advertisements were responsible for about 300 new subscribers. Would you recommend to Genie management that they purchase more Google ads?

Ans: Cost per subscriber = $9.99 per month

Cost per year = $9.99 *12 = $119.88

Total no. of subscribers = 300

Revenue for iGenie for 1 year after placing google advertisements= $119.88*300 = $35964

Cost for advertisements = $10000

Gross revenue after deducting expenses = $25964

Cost for CRM (customer relationship manager) = $10 per year

Total CRM cost = $10*300 = $3000

Revenue for iGenie after deducting CRM = $22964.

Retention rate = 20%

Discount rate for cost per subscriber = 8%

No.of subscribers for 2nd year = (300*20)/100 = 60

Cost per subscriber for month = (92*$9.99)/100 = $9.19

Cost per year = $9.19 *12 = $110.29

Revenue for 2nd year = $110.29 * 60 = $6617.38

Total CRM cost = 60*$10 =$600

Gross revenue = $6617.38 - $600 = $6017.38

Retention rate = 20%

Discount rate for cost per subscriber = 8%

No.of subscribers for 3rd year = (60*20)/100 = 12

Cost per subscriber for month = (92*$9.19)/100 = $8.45

Cost per year = $8.45 *12 = $101.46

Revenue for 3rd year = $101.46 * 12 = $1217.49

Total CRM cost = 12*$10 =$120

Gross revenue = $1217.49 - $120 = $1097.49

Retention rate = 20%

Discount rate for cost per subscriber = 8%

No.of subscribers for 4th   year = (12*20)/100 = 2

Cost per subscriber for month = (92*$8.45)/100 = $7.77

Cost per year = $7.77 *12 = $93.28

Revenue for 4th year = $93.28 * 2 = $187.57

Total CRM cost = 2*$10 =$20

Gross revenue = $187.57 - $20 = $167.57

For 5th year, there will not be any subscribers from the lot of 300, as after calculating retention rate number would be less than 1

So, total revenue generated after spending $10000 on Google ads is $22964+ $6017.38+$1097.49+$167.57 = $30246.44

Therefore, for every $10000 spent on Google ads, the overall net revenue earned is $30246.44

Hence, I would recommend that they purchase more google ads as it would generate more profit for iGenie.

  1. Suppose a newly-introduced loyalty program increases the number of customers that remained to 30%. Does this new data change your answer to 9.a?

Ans: If a new loyalty program is introduced which increases customers remained to 30%, the net revenue earned after spending $10000 on Google ads will increase. Also, the lot of 300 subscribers will earn revenue till 6th year, so more revenue can be generated.

Also, cost of new loyalty program should be taken into account. If the new loyalty program cost is minimal, this new data won’t change my answer. But, I would recommend iGenie can spend more on Google ads to increase their customer base.

  1. Do you have any hesitations or concerns about making recommendations to management based on your above estimate of customer lifetime value?

      Ans: I would have the following recommendations to management:

  • Introduce more customer loyalty programs such that retention rate would increase which increases the revenue earned for iGenie
  • Introduce life time packages with a minimum tenure of 10 years so that customer will sign after subscribing to the package
  • Include more features to the iGenie platform which would increase the subscriber base
  • Utilize budget allocated more judiciously and spend them in Google ads so that number of subscribers can be increased